Which of the following objects has the largest momentum?
(A) A bullet fired from a rifle
(B) A football quarterback running at top speed
(C) A horse walking at about 2 mileslhour
(D) An elephant standing still

Grade:upto college level

1 Answers

Deepak Patra
8 years ago
The correct option is (C).
Momentum (p) of an object is equal to the mass of the object (m) time’s speed of the object (v).
So, p = mv …… (1)
a) Let us consider the average mass of a bullet is 0.2 kg (m = 0.2 kg) and the average speed of bullet is 400 m/s (v = 400 m/s).
Using equation (1), the momentum of the bullet will be,
p = mv
= (0.2 kg) (400 m/s)
= 80 kg. m/s …… (2)
From equation (2) we observed that, the average momentum of the bullet will be 80 kg. m/s.
b) Let us consider the average mass of a football is 0.5 kg (m = 0.5 kg) and the top speed of football is 30 m/s (v = 30 m/s).
Using equation (1), the momentum of the football will be,
p = mv
= (0.5 kg) (30 m/s)
= 15 kg. m/s …… (3)
From equation (3) we observed that, the momentum of the football will be 11 kg. m/s.
c) Let us consider the average mass of a horse is 500 kg (m = 500 kg). The given speed of horse is 0.9 m/s (v = 2 miles/hour).
Using equation (1), the momentum of the horse will be,
p = mv
= (500 kg) (0.9 m/s)
= 450 kg. m/s …… (4)
From equation (4) we observed that, the average momentum of the horse will be 450 kg. m/s.
d) Since the elephant is still standing, therefore, the speed of the elephant would be zero (v=0). Using equation (1), the momentum of the elephant would be zero.
From the above observation we conclude that, a horse walking at about 2 miles/hour has the largest momentum. Thus option (C) is correct.